OpenStudy (anonymous):

add and write in simplist form: 5x+1/2x+1+3x+1/2x+1

OpenStudy (anonymous):

I'm thinking it's 2 + 9x....

OpenStudy (anonymous):

no i dont think so

OpenStudy (anonymous):

Let's add the numbers with the variables. 5x + 1/2x + 3x + 1/2x 1/2x + 1/2x = 1x 5x +1 x+3x = 9x Let's add the numbers without the variables. 1 + 1 = 2 Since you are just adding, it is 2x + 9.